# Are Ensembles Getting Better All the Time?

### v2

@@ D.1 Proof of Theorem 7
For all $i\geq 1$, let us set $Y_{i}\vcentcolon=X_{i}-\mu$. Note that, by definition, the $Y_{i}$s are i.i.d. and centered. We are going to apply two theorems of Petrov (1965) to the random variables $Y_{i}$s. Let us recall briefly Petrov’s results here. We start with some notation: $V(x)$ denotes the cumulative distribution function of $X_{1}$, $\mathfrak{U}$ denotes the points of increase of $V$ and $A$ its supremum, $\mathcal{B}^{+}$ denotes the set of non-negative values of $h$ such that $\int_{0}^{+\infty}\operatorname{e}^{hx}\mathrm{d}V(x)<+\infty$ and $B$ its supremum. Further, assuming $B>0$, for any $0<h<B$, we define
@@
where $u_{n}=o\left(1\right)$. Since $01/(1-\operatorname{e}^{-Hh})-1>$ 0, the previous reasoning applies, and $\operatorname{\mathbb{P}}\left(\overline{X}_{n}>\mu+\varepsilon\right)$ is also eventually monotonic. ∎
@@ D.2 Proof of Theorem 8
The result follows from this last display, since $\tau^{\top}a-\varphi(a)>0$. Indeed, $t\mapsto t^{\top}a-\varphi(t)$ is a strictly concave function, with maximum achieved at $t=\tau$ and value 00 at 00. All that is left to do is to prove Eq. (77). We use Theorem 1 in Joutard (2017). More precisely, we apply Theorem 1 to $Z_{n}=\overline{X}_{n}$, with rate $b_{n}=n$ and $H=0$. We have several assumptions to check.
@@ Appendix E When monotonicity does not hold
While innocent-looking, this last display has a complicated behavior. In particular, it is not monotonic (see Figure 7). Transforming slightly this example in order to satisfy the assumptions of Theorem 7 radically changes the behavior of $p_{n}$, as predicted. One possible way to do so is to add a small mass $\varepsilon/(2(\mu+\varepsilon)-1)$ at $\mu+\varepsilon$, while setting the probability to be equal to 00 and $1$ equal to $(2\mu-1)/(2(\mu+\varepsilon)-1)$. We refer to Figure 7 for an illustration.
@@ F.2 On the characteristic function of a random vector with non-zero density
First, let us suppose that there exists $t\neq 0$ such that $\operatorname{\mathbb{E}}\left[\mathrm{exp}\left(\mathbf{i}\langle t,X\rangle\right)\right]=1$. Then the expectation of the non-negative function $x\mapsto 1-\cos\langle t,x\rangle$ is exactly 00, meaning that $\langle t,x\rangle=0$ modulo $2\pi$ wherever $X$ has non-zero mass. Since $t\neq 0$, this implies that $X$ is supported in the collection of hyperplanes defined by

### v3

@@ D.1 Proof of Theorem 7
For all $i\geq 1$, let us set $Y_{i}\vcentcolon=X_{i}-\mu$. Note that, by definition, the $Y_{i}$s are i.i.d. and centered. We are going to apply two theorems of Petrov (1965) to the random variables $Y_{i}$s. Let us recall briefly Petrov’s results here. We start with some notation: $V(x)$ denotes the cumulative distribution function of $X_{1}$, $\mathfrak{U}$ denotes the points of increase of $V$ and $A$ its supremum, $\mathcal{B}^{+}$ denotes the set of non-negative values $h$ such that $\int_{0}^{+\infty}\operatorname{e}^{hx}\mathrm{d}V(x)<+\infty$ and $B$ its supremum. Further, assuming $B>0$, for any $0<h<B$, we define
@@
where $u_{n}=o\left(1\right)$. Since $1/(1-\operatorname{e}^{-Hh})-1>0$, the previous reasoning applies, and $\operatorname{\mathbb{P}}\left(\overline{X}_{n}>\mu+\varepsilon\right)$ is also eventually monotonic. ∎
@@ D.2 Proof of Theorem 8
The result follows from this last display, since $\tau^{\top}a-\varphi(a)>0$. Indeed, $t\mapsto t^{\top}a-\varphi(t)$ is a strictly concave function, with maximum achieved at $t=\tau$ and value 0 at 0. All that is left to do is to prove Eq. (53). We use Theorem 1 in Joutard (2017). More precisely, we apply Theorem 1 to $Z_{n}=\overline{X}_{n}$, with rate $b_{n}=n$ and $H=0$. We have several assumptions to check.
@@ Appendix E When monotonicity does not hold
While innocent-looking, this last display has a complicated behavior. In particular, it is not monotonic (see Figure 8). Transforming slightly this example in order to satisfy the assumptions of Theorem 7 radically changes the behavior of $p_{n}$, as predicted. One possible way to do so is to add a small mass $\varepsilon/(2(\mu+\varepsilon)-1)$ at $\mu+\varepsilon$, while setting the probability to be equal to 0 and $1$ equal to $(2\mu-1)/(2(\mu+\varepsilon)-1)$. We refer to Figure 8 for an illustration.
@@ F.2 On the characteristic function of a random vector with non-zero density
First, let us suppose that there exists $t\neq 0$ such that $\operatorname{\mathbb{E}}\left[\mathrm{exp}\left(\mathbf{i}\langle t,X\rangle\right)\right]=1$. Then the expectation of the non-negative function $x\mapsto 1-\cos\langle t,x\rangle$ is exactly 0, meaning that $\langle t,x\rangle=0$ modulo $2\pi$ wherever $X$ has non-zero mass. Since $t\neq 0$, this implies that $X$ is supported in the collection of hyperplanes defined by